# On Focusing Statistical Power for Searches and Measurements in Particle Physics

### v1

@@ I Introduction
A major theme in high-energy physics is the search for new phenomena and the desire to precisely measure known physics parameters. The design of such experiments can be expensive, and it can take decades to achieve the required sensitivity to weak or rare signals in an ocean of uninteresting backgrounds. A search in this context concludes when the existence of a particle has been confidently established after analyzing enough collisions, usually by rejecting the default hypothesis that the observed excess in a given sample would occur by chance, as with the Higgs boson discovery [1, 2]. Alternatively, the search may conclude by confidently ruling out the existence of a certain class of particles, as with several supersymmetry and dark matter searches [3, 4]. Similarly, a measurement is reported in the form of a confidence interval on a key property of the particle — a smaller interval means a more precise measurement.
Experimental particle physicists have almost universally relied on the generalized likelihood ratio test (LRT) for such studies [5, 6]. The LRT is optimal for simple-vs-simple hypothesis tests (Neyman-Pearson’s lemma, [7]), and also has convenient asymptotic properties [8, 9].
However, searches and measurements are typically performed as composite hypothesis tests. For instance, the measurement of a signal strength parameter for a potential new phenomenon or the mass of a known particle generally involves testing a default conjecture, such as $H_{0}:\mu=\mu_{0}$, against a composite alternative hypothesis, such as $H_{1}:\mu\neq\mu_{0}$, which encompasses a range of possible parameter values rather than a single one. In these cases, the Neyman–Pearson lemma no longer applies: the most powerful test may vary depending on the specific value $\mu_{1}$ within the alternative. In fact, for general settings, there is no single test that is simultaneously most powerful across all parameter values in the composite alternative. Effectively, this means that any choice of test statistic is implicitly a choice for how the statistical power of test varies over the parameter space; no test will be the most powerful everywhere.
@@
For our two case studies, we achieve significant gains using FTS in regimes with small sample sizes and low signal fractions, which is precisely where the most interesting physics can be found. Our method constructs confidence intervals via a Neyman construction, which is commonly implemented using a computationally expensive Feldman-Cousins technique [10] in particle physics. We also demonstrate a machine learning approach to rapidly perform Neyman construction [11, 12, 13] that provides valid confidence intervals for complex particle physics datasets. This technique applies to either test statistic and does not require large sample sizes, thereby enabling the construction of valid confidence intervals in a range of dark matter, neutrino or collider experiments where local asymptotics cannot be assumed [6, 14, 15, 16].
@@ II.1 How do we construct confidence intervals from tests?
Searches and measurements in particle physics are typically performed as hypothesis tests of the parameter $\mu$ of the likelihood $p(\mathcal{D}|\mu)$ implied by the physics processes under study for some parameter space $\Theta$ (See Appendix A for a detailed description of the experimental setup). For example, the generalized LRT rejects the null hypothesis $H_{0}:\mu=\mu_{0}$ (default conjecture on, e.g., the strength $\mu$ of a signal physics process) in favor of the alternative hypothesis $H_{1}:\mu\neq\mu_{0}$ when the generalized likelihood ratio statistic (LRS)
@@
is large or above some “critical value” $C_{\mu_{0}}$, which is chosen so that the proportion of wrongly rejected events (“false positives”) do not exceed some pre-set significance level $\alpha$. These hypothesis tests can then also be converted into an interval estimate of $\mu$ via the Neyman construction [17], where a $(1-\alpha)100$% confidence interval $I(\mathcal{D})$ of $\mu$ is defined by the set of all parameter values that passed the level-$\alpha$ tests:
@@ II.2 How do we cheaply compute critical values?
Data analyses in collider physics often rely on asymptotic properties (such as Wilks’ theorem [8, 9]) to cheaply estimate LRS confidence intervals, but asymptotics do not hold in small sample regimes. In this Letter, we propose a machine learning (ML) approach to quickly estimate how $C_{\mu_{0}}$ might vary with $\mu_{0}$; see Algorithm 1 in Appendix B for details of a procedure that performs a quantile regression of $T$ on $\mu_{0}$. An ML-based approach allows the data analyst to (i) take advantage of alternative test statistics (such as the proposed FTS defined below in Eq. 3), (ii) explore challenging small-sample regimes, and (iii) extend HEP analyses to higher-dimensional parameter spaces (such as in effective field theory) or to unbinned analyses, where a Monte Carlo (MC) approach (e.g., [6, Sec. 5.4]) is not computationally feasible.
@@ II.3 What test statistic can increase power in physics-motivated parameter regions?
[…]
which allows the data analyst to focus the power of a test or confidence procedure to specific regions of the parameter space, as defined by the focus or weight function $f$. In Bayesian testing, this statistic is known as a Bayes factor. Whereas the idea of using the Bayes factor as a frequentist test statistic is not new [18, 12], this work is the first to show that a test based on the Bayes Factor has the potential to improve upon likelihood ratio tests in particle physics experiments. For simplicity, we will in our experiments use a Gaussian focus function $f(\mu)$ with mean $m$ and standard deviation $s$, truncated so that $f(\mu)=0$ for $\mu\notin\Theta$. This choice may be guided by prior knowledge of $\mu_{0}$, for example, from an independent previous measurement or from theoretical considerations. The construction of an FTS confidence interval follows the procedure discussed above; see Figure 1(bottom).
@@ III Experiments and Results
We highlight two applications – a measurement of the coupling of Higgs bosons to tau leptons and a search for the existence of weakly interacting massive particles (WIMPs). The likelihoods are computed from bins of a histogram according to to the counting experiment convention as described in Appendix A. We compare LRS confidence intervals with FTS where we use truncated Gaussian focus functions of various widths (both narrow and wide choices), to demonstrate the robustness of our qualitative conclusions to these choices. To compute critical values, we adopt the procedure described in Section II.1.
Figure 1: Construction of confidence intervals. (Top) To construct a $(1-\alpha)100\%$ confidence interval using the LRT statistic (yellow; Eq. 1), we estimate the critical values $C_{\mu_{0}}$ (grey) via quantile regression and retain those $\mu_{0}$ values for which the statistic falls below the critical value. Shown are the critical values for $68\%$ (dark grey) and $95\%$ (light grey) confidence levels, with corresponding intervals displayed below the figure (dark and light yellow, respectively). (Bottom) The FTS statistic (blue; Eq. 3) for a focus function centered at $m=1$ and with width $s=1.2$ (dashed red) is compared against the critical values. Intervals at $68\%$ and $95\%$ confidence level are shown (dark and light blue, respectively). This figure corresponds to the Higgs measurement example using the Higgs mass observable, with the test statistics are evaluated on an Asimov data set [9]
@@ III.1 Measurement of a Higgs coupling via its signal strength
The goal of this case study is to measure the coupling of the Higgs boson to tau leptons [19]. This unknown quantity has a one-to-one correspondence with the signal strength in the Standard Model (SM) of particle physics and can therefore be posed as a signal strength measurement. The default hypothesis for the signal strength $\mu^{*}=1$ corresponds to the SM prediction. For this study, we let $\Theta=[0,10]$ and set the center of the focus function at $m=1$, the SM prediction. We define the wide focus width to be $s=2.4$, corresponding to the $2\sigma$ expected sensitivity of a prior measurement [20]; a narrow focus width of $s=1.2$ is also considered.
@@ III.1.1 Dataset
The dataset was simulated by the ATLAS experiment for the 2014 Higgs boson machine learning challenge (HiggsML) [21, 22]. It is comprised of a mixture of events obtained via several Higgs decay and background processes. For further details on the challenge, refer to [22].
In order to show the robustness of the analysis to our experimental design choices, we examine two natural choices of observables to build template histograms for the counting experiment setup. The first observable is the reconstructed Higgs mass; the second is the invariant mass of the hadronic tau and final-state lepton (“visible mass” for short). Both observables are known to discriminate between signal and background.
@@ III.1.2 Results
Estimation of critical values for both the LRS and FTS is accelerated by using quantile regression. With a low simulation budget of about $9{,}000$ pseudo-experiments (PEs), which for MC corresponds to 30 PEs per point times $300$ grid points, the mean-squared error (MSE) is improved by a factor of about 6 for LRS and 3 for FTS. See Figure 3 for an illustration of the relative speed-up.
@@ III.2 Search for WIMPs
Several experiments are dedicated to the search for dark matter candidates [23]. We design a data analysis scenario inspired by the search for WIMPs at the LZ experiment [24]. For our study, we construct a simplified dataset that mirrors the signal and background densities, as well as the sample size, of the recent LZ dark matter search [3]. The default hypothesis is that these dark matter candidates do not exist, corresponding to $\mu^{*}=0$; exclusion of this hypothesis would constitute a discovery. For this reason, our focus functions are centered at $m=0$. In this study, we let $\Theta=[0,40]$; the focus function is thus truncated below at 0.
@@ III.2.1 Dataset
We produce synthetic data mimicking the full dataset that passes selection cuts (as seen in [3, Fig. 3]) relating densities of background and 40 GeV WIMP models (see Appendix A.2 for details). For consistency with the previous example, we use two-dimensional histograms for density estimation, although these can readily be replaced with an analytical model for the density, when they are available in dark matter experiments [25]. The resulting 2D template histograms for the signal and background distributions are given in two observables, $S_{1}$ and $\log_{10}(S_{2})$, representing the light from prompt vacuum ultraviolet scintillation and delayed electroluminescence in the detector (see Fig. 4 in Appendix A.2). The simulated data are normalized such that when $\mu^{*}=1$, there are $1{,}200$ background events per $1$ signal event. The NEST simulator [26] was used to generate the template histograms.
@@ III.2.2 Results
In Figure 2b we find that FTS improves on LRS across the range of $0\leq\mu^{*}\leq 15$ with both wide ($s=6.0$) and narrow ($s=3.0$) focus functions. Table 1 (bottom two rows) shows that FTS with wide focus provides a median length $22\%$ shorter than LRS for a 68% confidence level in the setting where $\mu^{*}=0$. With narrow focus, the FTS intervals are $35\%$ shorter. As a result, we would expect a true negative search to conclude sooner with our method relative to the LRS-based method, on average. Furthermore, FTS maintains an advantage over LRS when $\mu^{*}\neq 0$. When $\mu^{*}=1$, the median $68\%$ confidence interval lengths for FTS are still $25\%$ shorter than LRS intervals when using wide focus and $34\%$ with narrow focus. If WIMPs exist, our method would find evidence for them sooner by excluding the default hypothesis using less data. See Appendix C for a discussion on setting upper and lower limits. As in the Higgs case, the absolute gains in median interval length are similar for 68% and 95% confidence levels. For $\mu^{*}=0$, there is 17% reduction for FTS-wide, 19% for FTS-narrow; and for $\mu^{*}=1$, a 16% reduction for FTS-wide and 16% for FTS-narrow.
@@
Figure 3: Critical values. (a) We compare the mean-squared error (MSE) in the critical value estimates obtained via MC (red) and quantile regression (QR, grey) for LRS on a grid of $300$ evaluation points. For a low simulation budget of $9{,}000$ pseudo-experiments (PEs), MC estimates show high MSE whereas QR estimates (using the same number of PEs) are comparable to MC estimates with $1.35$ million PEs. The right column confirms that low-budget QR yields accurate estimates across the parameter space, unlike low-budget MC. (b) For FTS, QR is again more efficient. With $9{,}000$ PEs, QR matches the performance of MC with about $30{,}000$ PEs.
@@ IV Discussion and Conclusion
The FTS framework gives the scientist control over the regions of maximum sensitivity, informed by independent previous measurements (as in our Higgs example) or by theory, without affecting the validity of the confidence intervals. As a theory-driven example, if an angular parameter is being measured, it may be sensible to concentrate power within $[0,2\pi]$ using, e.g., a von Mises (circular normal) distribution. The focus function allows the analyst to leverage prior knowledge about reasonable parameter regions within a fully frequentist analysis. However, the data under analysis should not be used to determine the focus. For example, one can explore the impact of different focus functions on the sensitivity using simulated samples over a range of true values of the parameter (see Figure 2) to provide guidelines, if the simulations did not use the observed data. The general rule is that a narrow focus leads to larger gains in power around the center of mass of the focus function, but one then pays a price (as compared to a wider focus) in terms of decreased power if the focus center of mass differs from the true value of the parameter $\mu^{*}$, as illustrated by the top row of Figure 2.
As an improvement to the core statistical methodology of the field, the FTS can enhance discovery potential and physics sensitivities across many experiments. Our case studies suggest that significant gains can be had when the signal-to-noise ratio is low and sample size is small, such as for dark matter searches, Higgs parameter measurements, and neutrino oscillation studies. Additionally, we provide a convenient ML tool that builds confidence intervals via a Neyman construction reliably and efficiently, with computational gains over traditional MC methods that are expected to scale with the dimensionality of the parameter space. The improved power of the FTS, enabled by the rapid building of the confidence intervals, can enhance the scientific discovery power of flagship measurements. Finally, our proposed method, which is here illustrated for histogram-based analyses, extends to unbinned, high-dimensional analyses with analytic expressions for the likelihood [25] as well to as any other parametric [27] or nonparametric likelihood estimators [28, 29, 30, 31, 32, 12].

### v2

@@ I Introduction
A major theme in high-energy physics is the search for new phenomena and the desire to precisely measure known physics parameters. The design of such experiments can be expensive, and it can take decades to achieve the required sensitivity to weak or rare signals in an ocean of uninteresting backgrounds. A search in this context concludes when the existence of a particle has been confidently established after analyzing enough collisions, usually by rejecting the default hypothesis that the observed excess in a given sample would occur by chance, as with the Higgs boson discovery [ATLAS:2012yve, CMS:2012qbp]. Alternatively, the search may conclude by confidently ruling out the existence of a certain class of particles, as with several supersymmetry and dark matter searches [LZ:2024zvo, Kwon:2025fgb]. Similarly, a measurement is reported in the form of a confidence interval on a key property of the particle — a smaller interval means a more precise measurement.
Experimental particle physicists have almost universally relied on the generalized likelihood ratio test (LRT) for such studies [cousins_lectures_2024, cranmer_practical_2015]. The LRT is optimal for simple-vs-simple hypothesis tests (Neyman-Pearson’s lemma, [neyman_ix_1933]), and also has convenient asymptotic properties [wilks_large-sample_1938, Cowan:2010js].
However, searches and measurements are typically performed as composite hypothesis tests. For instance, the measurement of a signal strength parameter for a potential new phenomenon or the mass of a known particle generally involves testing a default conjecture, such as $H_{0}:\mu=\mu_{0}$, against a composite alternative hypothesis, such as $H_{1}:\mu\neq\mu_{0}$, which encompasses a range of possible parameter values rather than a single one. In these cases, the Neyman–Pearson lemma no longer applies: the most powerful test may vary depending on the specific value $\mu_{1}$ within the alternative. In fact, for general settings, there is no single test that is simultaneously most powerful across all parameter values in the composite alternative. Effectively, this means that any choice of test statistic is implicitly a choice for how the statistical power of the test varies over the parameter space; no test will be the most powerful everywhere.
@@
For our two case studies, we achieve significant gains using FTS in regimes with small sample sizes and low signal fractions, which is precisely where the most interesting physics can be found. Our method constructs confidence intervals via a Neyman construction, which is commonly implemented using a computationally expensive Feldman-Cousins technique [Feldman:1997qc] in particle physics. We also demonstrate a machine learning approach to rapidly perform Neyman construction [pmlr-v119-dalmasso20a, dalmasso_likelihood-free_2024, masserano_simulator-based_2023] that provides valid confidence intervals for complex particle physics datasets. This technique applies to either test statistic and does not require large sample sizes, thereby enabling the construction of valid confidence intervals in a range of dark matter, neutrino or collider experiments where local asymptotics cannot be assumed [cranmer_practical_2015, NOvA:2017abs, ATLAS:2024jry, LZ:2025xkj].
@@ II.1 How do we construct confidence intervals from tests?
Searches and measurements in particle physics are typically performed as hypothesis tests on the parameter $\mu$ of the likelihood $p(\mathcal{D}|\mu)$ of datum $\mathcal{D}$ implied by the physics processes under study for some parameter space $\Theta$ (See Appendix A for a detailed description of the experimental setup). For example, the generalized LRT rejects the null hypothesis $H_{0}:\mu=\mu_{0}$ (default conjecture on, e.g., the strength $\mu$ of a signal physics process) in favor of the alternative hypothesis $H_{1}:\mu\neq\mu_{0}$ when the generalized likelihood ratio statistic (LRS)
@@
is large or above some “critical value” $C_{\mu_{0}}$, which is chosen so that the proportion of wrongly rejected events (“false positives”) do not exceed some pre-set significance level $\alpha$. These hypothesis tests can then also be converted into an interval estimate of $\mu$ via the Neyman construction [neyman_outline_1937], where a $(1-\alpha)100$% confidence interval $I(\mathcal{D})$ of $\mu$ is defined by the set of all parameter values that passed the level-$\alpha$ tests:
@@ II.2 How do we cheaply compute critical values?
Data analyses in collider physics often rely on asymptotic properties (such as Wilks’ theorem [wilks_large-sample_1938, Cowan:2010js]) to cheaply estimate LRS confidence intervals, but asymptotics do not hold in small sample regimes. In this Letter, we propose a machine learning (ML) approach to quickly estimate how $C_{\mu_{0}}$ might vary with $\mu_{0}$; see Algorithm 1 in Appendix B for details of a procedure that performs a quantile regression of $T$ on $\mu_{0}$. An ML-based approach allows the data analyst to (i) take advantage of alternative test statistics (such as the proposed FTS defined below in Eq. 3), (ii) explore challenging small-sample regimes, and (iii) extend HEP analyses to higher-dimensional parameter spaces (such as in effective field theory) or to unbinned analyses, where a Monte Carlo (MC) approach (e.g., [cranmer_practical_2015, Sec. 5.4]) is not computationally feasible.
@@ II.3 What test statistic can increase power in physics-motivated parameter regions?
$T_{f}(\mathcal{D};\mu_{0})=-2\log\left(\frac{p(\mathcal{D}|\mu_{0})}{\int_{\Theta}p(\mathcal{D}|\mu)f(\mu)d\mu}\right),$ (3)
which allows the data analyst to focus the power of a test or confidence procedure to specific regions of the parameter space, as defined by the focus or weight function $f$. In Bayesian testing, this statistic is known as a Bayes factor. Whereas the idea of using the Bayes factor as a frequentist test statistic is not new [good_bayesnon-bayes_1992, dalmasso_likelihood-free_2024, Fowlie03082023], this work is the first to demonstrate that a test based on the Bayes factor has the potential to improve upon likelihood ratio tests in various particle physics settings. For simplicity, we will in our experiments use a Gaussian focus function $f(\mu)$ with mean $m$ and standard deviation $s$, truncated so that $f(\mu)=0$ for $\mu\notin\Theta$. This choice may be guided by prior knowledge of $\mu_{0}$, for example, from an independent previous measurement or from theoretical considerations. The construction of an FTS confidence interval follows the procedure discussed above; see Figure 1 (bottom).
@@ III Experiments and Results
We highlight two applications – a measurement of the coupling of Higgs bosons to tau leptons and a search for the existence of weakly interacting massive particles (WIMPs). The likelihoods are computed from bins of a histogram according to the counting experiment convention as described in Appendix A. We compare LRS confidence intervals with FTS where we use truncated Gaussian focus functions of various widths (both narrow and wide choices), to demonstrate the robustness of our qualitative conclusions to these choices. To compute critical values, we adopt the procedure described in Section II.1.
Figure 1: Construction of confidence intervals. (Top) To construct a $(1-\alpha)100\%$ confidence interval using the LRS statistic (yellow; Eq. 1), we estimate the critical values $C_{\mu_{0}}$ (grey) via quantile regression and retain those $\mu_{0}$ values for which the statistic falls below the critical value. Shown are the critical values for $68\%$ (dark grey) and $95\%$ (light grey) confidence levels, with corresponding intervals displayed below the figure (dark and light yellow, respectively). (Bottom) The FTS statistic (blue; Eq. 3) for a focus function centered at $m=1$ and with width $s=1.2$ (dashed red) is compared against the critical values. Intervals at $68\%$ and $95\%$ confidence level are shown (dark and light blue, respectively). This figure corresponds to the Higgs measurement example using the Higgs mass observable, with the test statistics are evaluated on an Asimov data set [Cowan:2010js]
@@ III.1 Measurement of a Higgs coupling via its signal strength
The goal of this case study is to measure the coupling of the Higgs boson to tau leptons [Aad:2015vsa]. This unknown quantity has a one-to-one correspondence with the signal strength in the Standard Model (SM) of particle physics and can therefore be posed as a signal strength measurement. The default hypothesis for the true signal strength $\mu^{*}=1$ corresponds to the SM prediction. For this study, we let $\Theta=[0,10]$ to avoid right endpoint truncation bias in interval estimation, and we set the center of the focus function at $m=1$, the SM prediction. We define the wide focus width to be $s=2.4$, corresponding to the $2\sigma$ expected sensitivity of a prior measurement [ATLAS:2012gfw]; a narrow focus width of $s=1.2$ is also considered.
@@ III.1.1 Dataset
The dataset was simulated by the ATLAS experiment for the 2014 Higgs boson machine learning challenge (HiggsML) [HiggsMLOpenData, pmlr-v42-cowa14]. It is comprised of a mixture of events obtained via several Higgs decay and background processes. For further details on the challenge, refer to [pmlr-v42-cowa14].
In order to show the robustness of the analysis to our experimental design choices, we examine two natural choices of observables to build template histograms for the counting experiment setup. The first observable is the reconstructed mass of the Higgs boson candidate as estimated via phase space integration; the second is the invariant mass of the hadronic tau and final-state lepton (“visible mass” for short). Both observables are known to discriminate between signal and background with the reconstructed mass expected to provide stronger constraints.
@@ III.1.2 Results
Estimation of critical values for both the LRS and FTS is accelerated by using quantile regression. With a low simulation budget of about 9,0009{,}000 pseudo-experiments (PEs), which for MC corresponds to 30 PEs per point times $300$ grid points, the mean-squared error (MSE) is improved by a factor of about 6 for LRS and 3 for FTS. See Figure 3 for an illustration of the relative speed-up.
@@ III.2 Search for WIMPs
Several experiments are dedicated to the search for dark matter candidates [Bozorgnia:2024pwk]. We design a data analysis scenario inspired by the search for WIMPs at the LZ experiment [LZ:2019sgr]. For our study, we construct a simplified dataset that mimics the signal and background densities, as well as the sample size, of the recent LZ dark matter search [LZ:2024zvo]. The default hypothesis is that these dark matter candidates do not exist, corresponding to $\mu^{*}=0$; exclusion of this hypothesis would constitute a discovery. For this reason, our focus functions are centered at $m=0$. In this study, we let $\Theta=[0,40]$ to avoid upper endpoint truncation bias in interval estimation; the focus function is thus truncated below at 0.
@@ III.2.1 Dataset
We produce synthetic data mimicking the full dataset that passes selection cuts (as seen in [LZ:2024zvo, Fig. 3]) relating densities of background and 40 GeV WIMP models (see Appendix A.2 for details). For consistency with the previous example, we use two-dimensional histograms for density estimation, although these can readily be replaced with an analytical model for the density, when they are available in dark matter experiments [Aalbers:2020iej]. The resulting 2D template histograms for the signal and background distributions are given in two observables, $S_{1}$ and $\log_{10}(S_{2})$, representing the light from prompt vacuum ultraviolet scintillation and delayed electroluminescence in the detector (see Fig. 4 in Appendix A.2). The simulated data are normalized such that when $\mu^{*}=1$, there are 1,2001{,}200 background events per $1$ signal event. The NEST simulator [brodsky_nestcollaborationnest_2019] was used to generate the template histograms.
@@ III.2.2 Results
In Figure 2b we show median confidence interval length in the LZ-inspired experiment as a function of the true signal strength. We find that FTS improves on LRS across the range of $0\leq\mu^{*}\leq 15$ with both wide ($s=6.0$) and narrow ($s=3.0$) focus functions. Table 1 (bottom two rows) shows that FTS with wide focus provides a median length $22\%$ shorter than LRS for a 68% confidence level in the setting where $\mu^{*}=0$. With narrow focus, the FTS intervals are $35\%$ shorter. As a result, we would expect a true negative search to conclude sooner with our method relative to the LRS-based method, on average. Furthermore, FTS maintains an advantage over LRS when $\mu^{*}\neq 0$. When $\mu^{*}=1$, the median $68\%$ confidence interval lengths for FTS are still $25\%$ shorter than LRS intervals when using wide focus and $34\%$ with narrow focus. If WIMPs exist, our method would find evidence for them sooner by excluding the default hypothesis using less data. See Appendix C for a discussion on setting upper and lower limits. As in the Higgs case, the absolute gains in median interval length are similar for 68% and 95% confidence levels. For $\mu^{*}=0$, there is 17% reduction for FTS-wide, 19% for FTS-narrow; and for $\mu^{*}=1$, a 16% reduction for FTS-wide and 16% for FTS-narrow.
@@
Figure 3: Critical values. (a) We compare the mean-squared error (MSE) in the critical value estimates obtained via MC (red) and quantile regression (QR, grey) for LRS on a grid of 300300 evaluation points. For a low simulation budget of 9,0009{,}000 pseudo-experiments (PEs), MC estimates show high MSE whereas QR estimates (using the same number of PEs) are comparable to MC estimates with $1.35$ million PEs. The right column confirms that low-budget QR yields accurate estimates across the parameter space, unlike low-budget MC. (b) For FTS, QR is again more efficient. With 9,0009{,}000 PEs, QR matches the performance of MC with about 30,00030{,}000 PEs.
@@ IV Discussion and Conclusion
The FTS framework gives the scientist control over the regions of maximum sensitivity, whether chosen after optimization studies on independent simulations, informed by independent previous measurements (as in our Higgs example) or by theory, without affecting the validity of the confidence intervals. As a theory-driven example, if an angular parameter is being measured, it may be sensible to concentrate power within $0,2\pi]$ using, e.g., a von Mises (circular normal) distribution. The focus function allows the analyst to leverage prior knowledge about reasonable parameter regions within a fully frequentist analysis. However, the data under analysis should not be used to determine the focus. For example, one can explore the impact of different focus functions on the sensitivity using simulated samples over a range of true values of the parameter (see Figure 2) to provide guidelines, if the simulations did not use the observed data. The general rule is that a narrow focus leads to larger gains in power around the center of mass of the focus function, but one then pays a price (as compared to a wider focus) in terms of decreased power if the focus center of mass differs from the true value of the parameter $\mu^{*}$, as illustrated by the top row of Figure 2.
As an improvement to the core statistical methodology of the field, the FTS can enhance discovery potential and physics sensitivities across many experiments. Our case studies suggest that significant gains can be achieved when the signal-to-noise ratio is low and sample size is small, such as for dark matter searches, Higgs parameter measurements, and neutrino oscillation studies. Additionally, we provide a convenient ML tool that builds confidence intervals via a Neyman construction reliably and efficiently, with computational gains over traditional MC methods that are expected to scale with the dimensionality of the parameter space. The improved power of the FTS, enabled by the rapid building of the confidence intervals, can enhance the scientific discovery power of flagship measurements. Finally, our proposed method, which is here illustrated for histogram-based analyses, extends to unbinned, high-dimensional analyses with analytic expressions for the likelihood [Aalbers:2020iej] as well to as any other parametric [ATLAS:2018hxb] or nonparametric likelihood estimators [izbicki2014highdimensionaldensityratioestimation, Cranmer:2015bka, doi:10.1073/pnas.1912789117, ATLAS:2025clx, Ghosh:2025fma, dalmasso_likelihood-free_2024].